28 The following is a sample of 25 measurements. L02098 7 6 6 11 8 9 11 9 10 8 7 7 5 9 10 7 7 7 7 9 12 10 10 8 6 a. Compute , s2, and s for this sample b. Count the number of measurements in the intervals e ± s, x ± 2s, and x ± 3s. Express each count as a percentage of the total number of measurements. c. Compare the percentages found in part b with the percentages given by the empirical rule and Chebyshevs rule. d. Calculate the range and use it to obtain a rough approx- imation for s. Does the result compare favorably with the actual value for s found in part a?
Answer #1

(a)

Following table shows the calculations:

(x-mean)A2 1.5376 5.0176 5.0176 7.6176 0.0576 0.5776 7.6176 0.5776 3.0976 0.0576 1.5376 1.5376 10.4976 0.5776 3.0976 1.5376 1.5376 1.5376 1.5376 0.5776 14.1376 3.0976 3.0976 0.0576 5.0176 80.56 6 6 10 7 7 1) 10 7 12 10 10 6 Total 206

Sample size: n=25

Now,

Mean:

22. 206 = 8.24
The variance:

= 3.3567 n-1

Standard deviation:

2 (z-x) = 1.8321 ー

(b)

Following is the ordered data set:

10 10 10 10 12

First interval:

ř ± s = 8.24 ± 1.8321 = (6.4079. 10.0721)

Out of 25 data values, 18 lies in the above interval so the required percentage is

18 100%-72%

Second interval:

士2s 8.24 ± 2 . 1.8321 (4.5758. 11.9042)

Out of 25 data values, 24 lies in the above interval so the required percentage is

24 100% = 96%

Third interval:

z ± 3s 8.24 ± 3 . 1.8321 (2.7437. 13.7363)

Out of 25 data values, 25 lies in the above interval so the required percentage is

25 100%-100%

(c)

According to empirical rule 68% data values lies within one standard deviation of mean, 95% data values lies within two standard deviations of mean and 99.7% data values lies within three standard deviations of mean.

All the percentages of part b are greater than percentages given by empirical rule.

------------

Chebyshev's rule:

According to Chebyshev's rule at least 0% data values lies within one standard deviation of mean, at least 75% data values lies within two standard deviations of mean and at least 88.9% data values lies within three standard deviations of mean.

All the percentages of part b are as specified by part c.

(d)

The range is

Range = Maximum - Minimum = 12 - 5 = 7

According to range rule the standard deviation is approximately equal to one fourth of the range of the data so

7 s =-= 1.75

The standard deviation found in part (a) is greater than the above estimated standard deviation.